lexieec:

The function f(x) = 2x + 210 represents the number of calories burned when exercising, where x is the number of hours spent exercising. The function g(x) = 2x + 125 represents the calorie deficit that occurs when following a particular diet, where x is the number of hours spent exercising. What is (f + g)(1)? Explain. 339 calories burned while dieting for 1 hour. 339 calories burned while combining diet with 1 hour of exercise. 212 calorie calories burned when combining diet with 1 hour of exercise. 212 calories burned while exercising for 1 hour.

celticcat:

(f + g)(x) = 2x + 210 + 2x + 125 = 4x + 335 When x , the time, = 1 hour: (f + g)(1) = 4(1) + 335 = 339 cals. This is the calories burned when combining diet with 1 hour of exercise. The second option is the correct one.

jhonyy9:

nice work @celticcat but in future try collaborate with asker and please dont give directly answers just help with the way,step by step how he or she can solve the posted math problem